Refractive index of glass is 1.5. Find the wavelength of a beam of light with a frequency of 1014 Hz in glass. Assume velocity of light is 3 x 108 m/sec in vacuum.


1
Expert's answer
2021-01-06T13:57:25-0500

n = 1.5 (refractive index of a glass)

"\nu" = 1014Hz (frequancy)

c = 3 * 108 "\frac{m}{s}" (velocity of light) and "\lambda" (wavelength)

"\lambda" = "v * T = \frac{v}{\nu} = \frac{c}{n\nu} = \frac{3 * 10^8}{1.5*1014} =" 197239m